OpenStudy (anonymous):

Match each expression with its solution. Match Term Definition negative five−sixths minus one−half A) five−sixths negative two−thirds plus negative five−sixths B) negative one and one−half negative one−sixth plus two−thirds C) one−half one−half minus negative one−third D) negative one and one−third

OpenStudy (anonymous):

@natcatfinch121

OpenStudy (anonymous):

@harkirat

OpenStudy (math&ing001):

\[\frac{ 1 }{ 2 }-(-\frac{ 1 }{ 3 })=\frac{ 3 }{ 3*2 }+\frac{ 2 }{ 3*2 }=\frac{ 5 }{ 6 }\] It's pretty much the same for the others.